### v1

@@ 6.1 Inspection Schedules and Feedback Model
The model allows many inspection types: prefix-validity checks, milestone verification, partial correctness scores, or any measurement that is a function of the abstract state. The results below are information-theoretic and do not assume any particular learning algorithm.
@@ 6.2 Signal Decay to Intermediate Inspections
Proposition 6.3 implies that an inspection at time $u$ replaces the terminal attenuation factor $\mathrm{Attn}(t\to H)$ by the shorter-horizon factor $\mathrm{Attn}(t\to u)$.
@@
Corollary 6.4 makes the value of intermediate inspection precise: placing a checkpoint at time $u$ reduces the attenuation exponent from HtH-t to utu-t. The remaining question is how to place inspection times so that no step is too far from its next checkpoint.
@@ 6.3 Homogeneous Contraction: Minimax-Optimal Placement
The implication is that the detailed location of checkpoints is secondary; what matters is controlling the longest uninspected run. Combining Proposition 6.6 and Theorem 6.7 yields an explicit feasibility criterion.
@@
Example 6.10 (Semiconductor Fabrication)
A semiconductor process has $H=50$ stages with $\eta=0.85$. Terminal testing costs \$10/wafer; intermediate metrology costs \$50/inspection. With n=10,000n=10{,}000 wafers, $\Delta^{2}=0.2$, and target error $\epsilon=0.1$:
$H_{\mathrm{crit}}(\epsilon)=\frac{\ln(2000)-2\ln(0.9)}{\ln(1/0.85)}\approx 48.1<50.$
Since $H>H_{\mathrm{crit}}(\epsilon)$, outcome-only evaluation is insufficient. Corollary 6.9 gives $m\geq\lceil 50/48.1\rceil-1=1$. One inspection at the midpoint ($t_{1}=25$) reduces the maximal gap to 25, bringing all stages within the critical horizon. The inspection adds \$50/wafer but makes the study feasible; without it, the required sample size would exceed the budget.
@@ 6.4 Heterogeneous Contraction: Minimum Inspections via Information Distance
Because greedy stretches each segment to its feasibility limit, inspections concentrate automatically in regions where contraction is strongest.
When $\eta_{t}\equiv\eta$, we have $w_{t}\equiv\ln(1/\eta)$, so greedy produces segments of length $\lfloor\Gamma/\ln(1/\eta)\rfloor$, recovering uniform spacing (Theorem 6.7). When contraction varies, greedy produces non-uniform placement: segments spanning high-contraction steps are short; segments spanning low-contraction steps are long.
Example 6.14 (Service Operations)
A customer journey has $H=50$ touchpoints. Early touchpoints (awareness, consideration) have high variability ($\eta_{t}=0.6$ for $t\leq 10$), while later touchpoints are more deterministic ($\eta_{t}=0.95$ for $t>10$). With n=1,000n=1{,}000 customers, $\Delta^{2}=0.3$, and $\epsilon=0.1$, the feasibility threshold is $\Gamma=\ln(300)-2\ln(0.9)\approx 5.9$.
The 11 high-contraction steps contribute $W([0,11))=11\times\ln(1/0.6)\approx 5.6$, nearly exhausting the budget. The greedy algorithm extends through 5 additional low-contraction steps before exceeding $\Gamma$, placing an inspection at $t_{1}=16$ (since $W([0,16))\approx 5.9\leq\Gamma$ but $W([0,17))\approx 5.9>\Gamma$). The remaining 34 low-contraction steps contribute only $W([16,50))=34\times\ln(1/0.95)\approx 1.7<\Gamma$, fitting in one segment. The optimal schedule has one inspection, placed where contraction is strongest, demonstrating that non-uniform placement arises naturally when contraction varies.
@@ 6.5 Joint Sample-Inspection Budgeting
The preceding results characterize inspection placement in information units (maximal gap under homogeneous contraction, and information-distance segments under heterogeneous contraction). We now incorporate operational costs: in practice, the number of inspections is itself a design choice subject to a budget constraint. Let $c_{\mathrm{out}}$ denote the cost per trajectory for terminal evaluation and $c_{\mathrm{insp}}$ the cost per intermediate inspection per trajectory. Collecting $n$ trajectories with $m$ inspections each costs $B=n(c_{\mathrm{out}}+m\cdot c_{\mathrm{insp}})$.
@@
Proposition 6.15 reveals a tradeoff: increasing $m$ reduces the sample requirement exponentially, but raises per-trajectory cost linearly. A simple rule is to increase $m$ until the maximal gap $\lceil H/(m+1)\rceil$ falls below $H_{\mathrm{crit}}(\epsilon)$, at which point further inspections yield negligible sample savings at the target error level.
@@
Corollary 6.16 implies that inspection effort must grow nearly linearly with process depth to avoid exponential sample requirements. We summarize the preceding results in the following inspection design procedure.
Inspection Design Procedure 1. Estimate contraction: Bound the contraction coefficient $\eta$ (or $\{\eta_{t}\}$) using A/B comparisons, cohort tracking, or variance decay measurements. 2. Compute feasibility threshold: $\Gamma(n,\Delta^{2},\epsilon)=\ln(n\Delta^{2})-2\ln(1-\epsilon)$. Equivalently, under homogeneous contraction compute $H_{\mathrm{crit}}(\epsilon)=\Gamma/\ln(1/\eta)$. 3. Choose inspections for feasibility: Under homogeneous contraction, choose the smallest $m$ satisfying $m\geq\lceil H/H_{\mathrm{crit}}(\epsilon)\rceil-1$. Under heterogeneous contraction, choose the smallest $m$ produced by the greedy rule in Theorem 6.13. 4. Place inspections: Uniformly at $t_{i}=\lfloor iH/(m+1)\rfloor$ under homogeneous contraction (Theorem 6.7), or via greedy information-distance placement under heterogeneous contraction (Theorem 6.13). 5. Budget the study: Set $n$ and $m$ to satisfy $B=n(c_{\mathrm{out}}+mc_{\mathrm{insp}})$ and compare against the lower bound in Proposition 6.15.
The theoretical results yield three practical insights. First, intermediate inspection is often an investment in feasibility, not merely precision: a single well-placed checkpoint can make an otherwise impossible attribution problem tractable (Example 6.10). Second, uniform spacing provides a robust default when contraction rates are uncertain (Theorem 6.7); when stage-specific estimates are available, the greedy algorithm (Theorem 6.13) concentrates inspections where contraction is strongest. Third, inspection effort must scale with process depth: organizations extending their value chains cannot hold inspection budgets constant (Corollary 6.16).
@@ 7 Numerical Validation
We validate five core theoretical predictions using synthetic Markov chains with controlled parameters. For effective width and objective mismatch, we additionally demonstrate the predicted phenomena on LLM reasoning chains from GSM8K. Synthetic experiments allow exact control of the contraction coefficient $\eta$, providing clean verification of the bounds; LLM experiments confirm that the phenomena occur in practice. The experiments cover signal decay (Section 7.1), effective width under correlation (Section 7.2), optimal inspection design (Section 7.3), the critical horizon phase transition (Section 7.4), and the objective mismatch between additive and multiplicative rewards (Section 7.5). All code will be made available upon acceptance.
7.1 Signal Decay Bound
Theorem 3.9 establishes that $\chi^{2}$-divergence between outcome distributions decays exponentially with distance from outcome:
$\chi^{2}(P_{S_{H}}\|Q_{S_{H}})\;\leq\;\eta^{H-t}\cdot\chi^{2}(P_{S_{t}}\|Q_{S_{t}}).$
To test this bound, we construct synthetic Markov chains with 10 abstract states and exact contraction coefficient $\eta\in\{0.7,0.8,0.9,0.95\}$, using the kernel $K=\sqrt{\eta}\,I+(1-\sqrt{\eta})\,(1/10)\,\mathbf{1}\mathbf{1}^{\top}$, a convex combination of the identity and the uniform-mixing matrix whose $\chi^{2}$-contraction coefficient is exactly $\eta$. For each step $t$, we set $P_{t}=\delta_{0}$ (point mass on state 0) and $Q_{t}=\pi$ (the uniform stationary distribution), giving $\chi^{2}(P_{t}\|Q_{t})=9$. We propagate both distributions to the terminal state through the contractive kernel and measure the resulting $\chi^{2}$-divergence at the state distribution level.
Figure 7.1 plots the measured divergence against distance to outcome (Ht)(H-t) on a logarithmic scale for $H=40$, showing all four values of $\eta$. The dashed lines show the theoretical prediction $\eta^{H-t}\cdot\Delta^{2}$ with $\Delta^{2}=9$. Because the reference distribution $Q_{t}$ equals the stationary distribution $\pi$, the bound is achieved with equality: measured divergences lie exactly on the theoretical lines. The measurements follow perfect linear trends on the log scale with distinct slopes for each $\eta$, confirming exponential decay at the predicted rates. Results for other horizons ($H\in\{10,20\}$) are identical in pattern.
@@ 7.1 Signal Decay Bound
Validation of the Signal Decay Bound (Theorem 3.9). Measured $\chi^{2}$-divergences (dots) match the theoretical prediction $\eta^{H-t}\cdot\Delta^{2}$ (dashed lines) exactly for all four contraction rates, confirming exponential decay with rate $\eta$. Setup: 10-state Markov chains with $H=40$ and $\Delta^{2}=9$.
7.2 Effective Width Under Correlation
Theorem 4.4 shows that positively correlated rollouts yield diminished variance reduction:
$W_{\mathrm{eff}}\;=\;\frac{W}{1+(W-1)\rho},$
where $\rho$ is the pairwise correlation between rollout outcomes. As $W\to\infty$, effective width saturates at $1/\rho$, meaning that beyond a certain point, additional rollouts provide negligible benefit.
We test this prediction in two settings. In synthetic experiments, we induce correlation $\rho=0.15$ by adding a shared noise component across rollouts, then measure variance reduction at $W\in\{1,4,16,64,256\}$. In LLM experiments, we generate $W\in\{1,4,16\}$ rollouts per GSM8K problem and measure the empirical pairwise correlation ($\bar{\rho}\approx 0.49$, averaged across $W\in\{4,16\}$) and variance reduction.
Figure 7.2 shows the results. In both panels, measured effective width (red dots) closely tracks the theoretical prediction (blue curve). The synthetic experiment confirms saturation: at $W=256$ rollouts, effective width reaches only $6.6$, compared to the theoretical maximum of $1/\rho=6.7$. The LLM experiment shows even stronger saturation due to higher correlation: $W=16$ nominal rollouts yield effective width of only $2.2$.
[…]
Validation of the Effective Width Formula (Theorem 4.4). Measured effective width (dots) matches the theoretical prediction $W/(1+(W-1)\rho)$ (curve) and saturates at $1/\rho$. (a) Synthetic with $\rho=0.15$. (b) LLM with $\rho\approx 0.49$.
7.3 Optimal Inspection Design
Theorem 6.7 establishes that under homogeneous contraction, uniformly spaced checkpoints minimize the worst-case attribution error across all steps. The intuition is that any gap between checkpoints allows signal to decay; uniform spacing ensures no gap is larger than necessary.
We test this with $H=20$ steps, contraction coefficient $\eta=0.9$, and $m=3$ intermediate checkpoints, comparing four schedules: uniform placement at $\{5,10,15\}$, front-loaded at $\{2,4,6\}$, back-loaded at $\{14,16,18\}$, and random at $\{2,13,14\}$. For each schedule, we compute the worst-case attribution error across all 20 steps.
Figure 7.3 shows the results. Uniform spacing achieves worst-case error of 0.41, while the alternatives achieve 0.69 (random) to 0.77 (front- and back-loaded). This represents a 40–47% reduction in worst-case error, confirming that uniform spacing is not merely a reasonable heuristic but the optimal choice under the conditions of the theorem.
[…]
Validation of Optimal Inspection Design (Theorem 6.7). Uniform checkpoint spacing (green) achieves 40–47% lower worst-case error than front-loaded, back-loaded, or random schedules. Checkpoint positions shown above each bar.
7.4 Critical Horizon Phase Transition
Corollary 3.10 predicts a sharp phase transition: steps within $H_{\mathrm{crit}}\approx\ln(n\Delta^{2})/\ln(1/\eta)$ of the outcome can be reliably attributed, while steps beyond this horizon cannot. We test this prediction by measuring attribution accuracy as a function of distance from outcome.
We use $H=40$ steps, 10 abstract states, n=1,000n=1{,}000 samples per hypothesis, and $\Delta^{2}=9$. We test two contraction rates: $\eta=0.7$ (predicted $H_{\mathrm{crit}}\approx 26$) and $\eta=0.8$ (predicted $H_{\mathrm{crit}}\approx 41$). For each step position, we generate outcome samples under two initial distributions and test whether the step’s identity can be recovered.
Figure 7.4 shows the results. Attribution accuracy for $\eta=0.7$ drops from 89% at distance 1 to chance level (50%) around distance 23, slightly before the theoretical $H_{\mathrm{crit}}\approx 26$. For $\eta=0.8$, the slower decay rate places $H_{\mathrm{crit}}\approx 41$ beyond the maximum horizon $H=40$, so accuracy remains above chance throughout but declines steadily. The critical horizon formula provides a tight upper bound on the reach of outcome-based credit assignment: accuracy degrades continuously with distance and reaches chance level at or before $H_{\mathrm{crit}}$, confirming that steps beyond the critical horizon cannot be reliably evaluated. The gap between the empirical transition and the theoretical bound reflects the difference between the simple classifier used here and the information-theoretically optimal test.
[…]
Validation of the Critical Horizon (Corollary 3.10). Attribution accuracy drops to chance (50%) at or before the predicted $H_{\mathrm{crit}}$ (dashed lines) for both contraction rates. For $\eta=0.7$, $H_{\mathrm{crit}}\approx 26$; for $\eta=0.8$, $H_{\mathrm{crit}}\approx 41$, which exceeds the horizon $H=40$, so accuracy remains above chance throughout.
7.5 Objective Mismatch
Definition 5.1 identifies a fundamental gap between step-quality ($J_{\mathrm{add}}$) and validity ($J_{\mathrm{mult}}$) objectives: a chain can have most steps correct yet still produce a wrong answer. We test whether this “mostly correct fallacy” occurs in practice by evaluating 50 GSM8K reasoning chains, scoring each step for correctness and comparing additive rewards (fraction of steps rated correct) against a binary validity indicator (1 if the final answer is correct, 0 otherwise).
Figure 7.5(a) plots each chain’s additive reward (x-axis) against its validity indicator (y-axis: 1 for correct, 0 for wrong). Dots along the top are chains that produce a correct final answer; dots along the bottom are chains with a wrong final answer. The shaded region marks the “mostly correct but wrong” zone: additive reward $\geq 0.80$ yet validity =0=0. Eight chains (16%) fall in this zone. Six of these have every step rated correct (additive reward =1.0=1.0) but still produce a wrong final answer, because the step-level heuristic evaluates numerical consistency of intermediate calculations but cannot catch errors in problem setup or final answer extraction. Panel (b) shows the distribution of additive rewards separately for correct-answer and wrong-answer chains; the dashed line marks the $0.80$ threshold. Both correct-answer and wrong-answer chains concentrate at high additive reward, confirming that step-level scores do not reliably separate valid from invalid reasoning. These chains span 23–48 reasoning steps, so the phenomenon is not an artifact of short chains. An additive objective would assign near-maximal reward to all eight, reinforcing the reasoning patterns that produce wrong answers; the validity indicator correctly identifies them as failures.
[…]
Validation of the Objective Mismatch (Definition 5.1). (a) Additive vs. multiplicative reward for 50 GSM8K chains. The shaded region highlights 8 chains (16%) that are “mostly correct” (high additive reward) but wrong (multiplicative reward =0=0). (b) Distribution of step correctness, showing that wrong-answer chains can have nearly all steps correct.
@@ 8 Conclusion
This paper establishes information-theoretic limits on credit assignment in deep sequential processes, revealing a phase transition that constrains both inspection system design and reasoning model training.
@@
The Signal Decay Bound (Theorem 3.9) shows that information about step $t$ decays as $\eta^{H-t}$, defining a critical horizon:
$H_{\mathrm{crit}}=\frac{\ln(n\Delta^{2})}{\ln(1/\eta)}.$
Steps within $H_{\mathrm{crit}}$ of the outcome can be evaluated from endpoint data; steps beyond cannot, regardless of sample size or algorithmic sophistication.
@@
Several limitations merit acknowledgment. The contraction coefficient $\eta$ is typically unknown and must be estimated empirically. The semantic state abstraction is posited conceptually but not constructively defined; developing practical methods to identify semantic states from token sequences remains open. Binary validity is a simplification; real reasoning admits degrees of correctness, partial credit, and recoverable errors.
Extending the framework in three directions would strengthen its practical impact: (1) graded validity models that distinguish fatal errors from recoverable mistakes; (2) empirical estimation procedures for $\eta$ and $\Delta^{2}$ from observational data; and (3) adaptive inspection schemes that learn contraction parameters online while optimizing checkpoint placement.
In summary.
Our results quantify exactly when outcome supervision suffices and when intermediate feedback is provably necessary:
Condition Implication
$H<H_{\mathrm{crit}}$ Outcome supervision suffices
$H>H_{\mathrm{crit}}$ Process supervision required
$\rho>0$ Effective width capped at $1/\rho$
Additive objective Optimizes wrong quantity for validity
Homogeneous contraction Uniform checkpoint spacing is minimax-optimal

### v2

@@ 6.1 Inspection Schedules and Feedback Model
The model allows many inspection types: prefix-validity checks, milestone verification, partial correctness scores, or any measurement that is a function of the abstract state. Remark 14.2 in Appendix 14 extends the model to noisy inspections with false positives and negatives, adding an observation-channel penalty to the segment attenuation budget. The results below are information-theoretic and do not assume any particular learning algorithm.
@@ 6.2 Signal Decay to Intermediate Inspections
Proposition 6.3 implies that an inspection at time $u$ replaces the terminal attenuation factor $\mathrm{Attn}(t\to H)$ by the shorter-horizon factor $\mathrm{Attn}(t\to u)$. The following corollary makes the value of intermediate inspection precise: placing a checkpoint at time $u$ reduces the attenuation exponent from HtH-t to utu-t. The remaining question is how to place inspection times so that no step is too far from its next checkpoint.
@@ 6.3 Homogeneous Contraction: Minimax-Optimal Placement
While uniform spacing is a classical heuristic for worst-case gap objectives, our contribution is identifying the correct design objective, namely the gap in SDPI attenuation units, and linking it to sample complexity via the Signal Decay Bound. The implication is that the detailed location of checkpoints is secondary; what matters is controlling the longest uninspected run. Combining Proposition 6.6 and Theorem 6.7 yields an explicit feasibility criterion.
@@
Example 15.2 in Appendix 15 applies Corollary 6.9 to a 50-stage semiconductor fabrication process, showing that a single midpoint inspection suffices to bring all stages within the critical horizon.
@@ 6.4 Heterogeneous Contraction: Minimum Inspections via Information Distance
Greedy segmentation is standard for interval-packing problems; what the SDPI analysis adds is the metric of cumulative information distance $W([t_{i},t_{i+1}))=\sum_{t}\ln(1/\eta_{t})$ rather than step count, under which feasibility is defined, so inspections concentrate automatically where contraction is strongest. When $\eta_{t}\equiv\eta$, we have $w_{t}\equiv\ln(1/\eta)$, so greedy produces segments of length $\lfloor\Gamma/\ln(1/\eta)\rfloor$, recovering uniform spacing (Theorem 6.7). When contraction varies, greedy produces non-uniform placement: segments spanning high-contraction steps are short; segments spanning low-contraction steps are long. Example 15.3 in Appendix 15 applies Theorem 6.12 to a 50-touchpoint customer journey with depth-dependent contraction, showing that the greedy algorithm places its single inspection in the high-contraction early stages.
@@ 6.5 Joint Sample-Inspection Budgeting
Inspection frequency trades off per-trajectory cost against the exponential sample savings from shortening information paths. Let $c_{\mathrm{out}}$ denote the per-trajectory cost of terminal evaluation and $c_{\mathrm{insp}}$ the per-trajectory cost per intermediate inspection. Collecting $n$ trajectories with $m$ inspections each has total budget $B=n(c_{\mathrm{out}}+mc_{\mathrm{insp}})$.
@@
Increasing $m$ reduces the sample requirement exponentially but raises per-trajectory cost linearly. A practical rule is to choose the smallest $m$ that makes the uniform segment length $\lceil H/(m+1)\rceil$ fall within $H_{\mathrm{crit}}(\epsilon)$ (Corollary 6.9).
@@
Corollary 6.14 implies that inspection effort must grow nearly linearly with process depth to avoid exponential sample requirements. An implementable inspection design procedure and detailed cost–sample tradeoff analysis are provided in the Electronic Companion (Appendix 16).
@@ 7 Numerical Validation
We validate the paper’s five core predictions through synthetic Markov chains with controlled parameters and, where applicable, LLM reasoning chains on GSM8K (a grade-school math benchmark). Across all experiments, observed behavior closely matches the theoretical bounds, confirming that the closed-form results provide reliable quantitative guidance in practice. Table 7 summarizes the headline findings; full experimental setups, parameter configurations, and all figures are provided in the Electronic Companion (Appendix 17).
@@
Headline numerical findings. Full setups and plots in Appendix 17. Prediction Test bed Key finding Signal decay (Thm. 3.8) 10-state synthetic chains, $\eta\in\{0.7,0.8,0.9,0.95\}$ Measured $\chi^{2}$-divergences match $\eta^{H-t}\Delta^{2}$ exactly ($R^{2}>0.999$). Effective width (Thm. 4.3) Synthetic ($\rho=0.15$); LLM rollouts on GSM8K ($\rho\approx 0.49$) $W_{\mathrm{eff}}$ saturates at $6.6$ vs. theoretical $1/\rho=6.7$ (synthetic); at $2.2$ vs. $2.0$ (LLM). Uniform inspection (Thm. 6.7) $H=20$, $\eta=0.9$, $m=3$ checkpoints Uniform spacing achieves 40–47% lower worst-case error than front-loaded, back-loaded, or random schedules. Critical horizon (Cor. 3.9) Synthetic, $\eta\in\{0.7,0.8\}$, n=1,000n=1{,}000 Attribution accuracy drops to chance ($50\%$) at or before the predicted $H_{\mathrm{crit}}$. Objective mismatch (Def. 5.1) 50 LLM reasoning chains on GSM8K 16% of chains are “mostly correct” (additive reward $\geq 0.80$) yet produce wrong answers.
@@ 8 Conclusion
This paper establishes information-theoretic limits on credit assignment in deep sequential processes, revealing a sharp sample-complexity threshold that constrains both inspection system design and reasoning model training.
@@
The Signal Decay Bound (Theorem 3.8) shows that information about step $t$ decays as $\eta^{H-t}$, defining a critical horizon $H_{\mathrm{crit}}\approx\ln(n\Delta^{2})/\ln(1/\eta)$ (up to the $\epsilon$-correction in Corollary 3.9). Steps within $H_{\mathrm{crit}}$ of the outcome can be evaluated from endpoint data; steps beyond require exponentially many samples in the distance HtH-t.
@@
Key extensions include data-driven estimation of $\eta$ and $\Delta^{2}$, learning the abstraction $\phi$ from observations, and moving from binary to graded validity.